# ACC-Debate: An Actor-Critic Approach to Multi-Agent Debate

### v1

@@ 4.3 Off-Policy Trajectory Generation
Guided-Debate To address these limitations and improve efficiency, we propose guided-debate trajectories, which steer the debate procedure in two opposing directions: one towards, and another away from, the correct answer. By comparing these guided trajectories with the natural debate trajectory, we can assess the relative goodness of each trajectory using an estimation of the reward structure $r$. To guide debate trajectories, we take inspiration from Zelikman et al. (2022) and use prompt modifications to provide additional information to models such that their responses support either a correct or incorrect answer.
@@
More specifically, for task $x$ with answer $y$, let $\mathbf{z}^{(t-1)}=(z_{a}^{(t-1)},z_{c}^{(t-1)})$ be the models’ responses at time $t-1$. Let $(z_{a}^{(t)},z_{c}^{(t)})$ be the models’ natural responses (i.e., without guidance), let $(z_{y,a}^{(t)},z_{y,c}^{(t)})$ and $(z_{!y,a}^{(t)},z_{!y,c}^{(t)})$ be the models responses when guided towards, and away from, supporting answer $y$ respectively. Thus, each guided response is an off-policy generation. In practice, we want guided responses to be different enough from natural responses so that learning the guided responses results in consequential changes to the model, but not so different that they are challenging to learn; we find that prompt modification is an effective tool for striking this balance. To guide the generations of $(z_{y,a}^{(t)},z_{y,c}^{(t)})$ and $(z_{!y,a}^{(t)},z_{!y,c}^{(t)})$, we will simply provide a correct and wrong target answer in the prompt, respectively - see “Guided Debate Prompts” in Section C of the supplement for further details.
@@
The terms $\Delta_{y}$ and $\Delta_{!y}$ give the expected accuracy difference if at round $t$ the actor had given response $z_{y,a}^{(t)}$ (or $z_{!y,a}^{(t)}$) instead of response $z_{a}^{(t)}$. Large $\Delta_{y}$ indicates that a one-response difference during the debate was sufficient to push the procedure toward the correct answer. Such responses would be desirable for the model to learn. On the other hand, large values of $\Delta_{!y}$ indicate that a one-response difference easily causes the models to converge to the incorrect answer; this indicates that the debate procedure is particularly fragile at timestep $t$.
With these observations in hand, we use $\Delta_{y}$ and $\Delta_{!y}$ to define positive and negative examples, in particular for a threshold $\varepsilon$, (if neither value is above the threshold, then the example is thrown out).
$(z_{+}^{(t)},z_{-}^{(t)})=\begin{cases}(z_{y}^{(t)},z^{(t)})&\text{if}\varepsilon\leq\Delta_{y}=r(z_{y,a}^{(t)},x,y)-r(z_{a}^{(t)},x,y)\\ (z^{(t)},z_{!y}^{(t)})&\text{if}\varepsilon\leq\Delta_{!y}=r(z_{a}^{(t)},x,y)-r(z_{!y,a}^{(t)},x,y)\end{cases}$ (4)
@@
Under Eq. 4, a positive example $z_{+}^{(t)}$ can be interpreted as a guided response $z_{y}^{(t)}$ which increased the probability of debate converging to the correct answer by at least $\varepsilon$, when compared with the natural response $z^{(t)}$. Similarly, a negative example $z_{-}^{(t)}$ is a guided response $z_{y!}^{(t)}$ which decreased the probability of debate converging to the answer by at least $\varepsilon$.
Now that we have a procedure for generating high-quality training examples consisting of positive and negative pairs, we next discuss how to use those positive and negative pairs to train both models.
Preference Optimization
@@
Hence, given a preference dataset of positive and negative examples for both the actor and critic model, of the from $z_{-}^{(t)},z_{+}^{(t)}$, the DPO loss is defined as,
$\mathcal{L}_{\text{DPO}}=\sum_{t=0}^{T}\mathbb{E}_{(x,y,z_{-}^{(t)},z_{+}^{(t)})\sim D}\bigg{[}\log\sigma\bigg{(}\frac{\pi_{\theta}\big{(}z_{+}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}{\pi_{\text{ref}}\big{(}z_{+}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}-\frac{\pi_{\theta}\big{(}z_{-}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}{\pi_{\text{ref}}\big{(}z_{-}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}\bigg{)}\bigg{]}$
Where $\pi_{\theta}$ is the policy induced by parameters $\theta_{a}$ or $\theta_{c}$ and $\mathbf{z}^{(t-1)}$ are the model’s responses at the previous round (i.e., the responses before giving either response $z_{-}^{(t)}$ or $z_{+}^{(t)}$).
@@
This loss implicitly optimizes the reward $r(z^{t},x,y)$, which is equivalent to final debate accuracy (i.e., the quantity being maximized by Eq. 1). By summing across all rounds, we implicitly maximize the probability that each round $t$ yields a response $z^{t}$ which causes debate to converge to the correct answer at time $T$.
@@ 5 Experiments
To evaluate the efficacy of ACC-Debate  we make use of 5 standard benchmark tasks: BoolQ Clark et al. (2019) $\sim$12k yes-no reading comprehension questions, MMLU Hendrycks et al. (2020) $\sim$15k multiple choice questions covering a wide array of subjects and difficulty, BBH Suzgun et al. (2022) $\sim$5k mixed-type questions SCIQ Welbl et al. (2017) $\sim$13k multiple-choice science questions, ARC Chollet (2019) $\sim$7k multiple-choice reasoning-based questions.
@@
We compare ACC-Debate to several debate baselines. For inference-based methods, we compare to Society of Minds SoM Du et al. (2023), Persona Chan et al. (2023). For training-based methods, we compare to supervised fine-tuning SFT Radford (2018), DebateTune Li et al. (2024b), DebateGPT Subramaniam et al. (2024). We perform experiments using three different base models: Llama-3-8B-Instruct Llama-3 Dubey et al. (2024), Mistral-7B-Instruct Mistral Jiang et al. (2023), and Gemma-2-2B-Instruct Gemma-2 Team et al. (2024).
5.1 ACC-Debate Performance
@@
We begin by examining the performance of our method ACC-Debate (a single round of training) and ACC-Debate+ (two rounds of training) In table 1, we see the average accuracy of each method after five rounds of debate. Our method attains superior performance compared with baseline methods in most cases. The high efficacy of ACC-Debate relative to the baselines indicates that in most cases, only a single round of training is necessary to produce a high-quality debate team. It is worth noting that in some cases, further training rounds may decrease performance (i.e., ACC-Debate+ can have worse performance than ACC-Debate). As such, we recommend maintaining a hold-out set of tasks to determine whether further training degrades performance.
@@
Llama-3             (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Debate ACC-Debate+ BoolQ .$812_{\pm.01}.$$811_{\pm.007}.$$781_{\pm.002}.$$775_{\pm.033}.$$798_{\pm.006}.$$815_{\pm.005}.$$887_{\pm.005}.$$\mathbf{.894}_{\pm.003}$ MMLU .$62_{\pm.004}.$$635_{\pm.004}.$$639_{\pm.004}.$$63_{\pm.004}.$$642_{\pm.005}.$$654_{\pm.005}.$$644_{\pm.01}.$$\mathbf{.683}_{\pm.012}$ BBH .$508_{\pm.003}.$$514_{\pm.005}.$$509_{\pm.013}.$$508_{\pm.005}.$$552_{\pm.006}.$$551_{\pm.008}.$$\mathbf{.593}_{\pm.006}$$574_{\pm.003}.$ SCIQ .$925_{\pm.002}.$$923_{\pm.002}.$$925_{\pm.004}.$$924_{\pm.004}.$$925_{\pm.003}.$$932_{\pm.001}.$$\mathbf{.952}_{\pm.0}$$948_{\pm.003}.$ ARC .$874_{\pm.001}.$$874_{\pm.001}.$$87_{\pm.003}.$$871_{\pm.002}.$$879_{\pm.004}.$$876_{\pm.002}.$$\mathbf{.881}_{\pm.004}$$869_{\pm.002}.$ Mistral               (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Debate ACC-Debate+ BoolQ .$801_{\pm.005}.$$798_{\pm.004}.$$831_{\pm.003}.$$83_{\pm.003}.$$84_{\pm.003}.$$848_{\pm.002}.$$877_{\pm.002}.$$\mathbf{.893}_{\pm.002}$ MMLU .$57_{\pm.003}.$$562_{\pm.005}.$$574_{\pm.002}.$$562_{\pm.005}.$$594_{\pm.004}.$$577_{\pm.002}.$$61_{\pm.005}.$$\mathbf{.672}_{\pm.004}$ BBH .$428_{\pm.002}.$$462_{\pm.003}.$$465_{\pm.011}.$$456_{\pm.005}.$$439_{\pm.006}.$$48_{\pm.012}.$$519_{\pm.009}.$$\mathbf{.601}_{\pm.004}$ SCIQ .$856_{\pm.002}.$$856_{\pm.002}.$$86_{\pm.002}.$$863_{\pm.003}.$$858_{\pm.004}.$$871_{\pm.003}.$$902_{\pm.005}.$$\mathbf{.905}_{\pm.002}$ ARC .$824_{\pm.001}.$$823_{\pm.002}.$$827_{\pm.001}.$$834_{\pm.0}.$$825_{\pm.003}.$$822_{\pm.002}.$$843_{\pm.003}.$$\mathbf{.856}_{\pm.003}$ Gemma-2              (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Debate ACC-Debate+ BoolQ .$75_{\pm.011}.$$759_{\pm.004}.$$716_{\pm.015}.$$767_{\pm.003}.$$783_{\pm.011}.$$812_{\pm.003}.$$84_{\pm.005}.$$\mathbf{.845}_{\pm.005}$ MMLU .$58_{\pm.002}.$$578_{\pm.002}.$$577_{\pm.002}.$$578_{\pm.001}.$$579_{\pm.002}.$$\mathbf{.582}_{\pm.002}$$51_{\pm.016}.$$555_{\pm.003}.$ BBH .$454_{\pm.007}.$$449_{\pm.01}.$$447_{\pm.006}.$$447_{\pm.007}.$$498_{\pm.006}.$$491_{\pm.01}.$$\mathbf{.513}_{\pm.006}$$475_{\pm.008}.$ SCIQ .$903_{\pm.002}.$$903_{\pm.002}.$$908_{\pm.003}.$$903_{\pm.001}.$$913_{\pm.002}.$$914_{\pm.002}.$$\mathbf{.918}_{\pm.003}$$909_{\pm.003}.$ ARC .$841_{\pm.003}.$$843_{\pm.005}.$$847_{\pm.003}.$$847_{\pm.003}.$$848_{\pm.002}.$$851_{\pm.003}.$$\mathbf{.852}_{\pm.003}$$849_{\pm.002}.$
@@
Table 1: Average accuracy (with 95% confidence intervals) after $5$ rounds of debate.
5.2 Performance Increase of Debate
@@
Figure 2: Percent improvement in accuracy after five rounds of debate, compared to a single round. Percent improvement (Eq. 5) for each method is averaged across all five datasets.
@@
Average Improvement As noted in Du et al. (2023), the key mechanism behind the success of multi-agent debate (or any of its many variants) is that discussion over multiple rounds allows the models to iteratively refine their answers. Thus, a natural question is: how much does accuracy improve from the initial round $t=0$ to the final round $t=T$? To measure this, we look at the percent improvement in model accuracy from round $t=0$ to round $t=4$ calculated as,
@@
In Figure 2 we see the average percent improvement for each method, averaged across all 5 datasets. For each of the three base models, ACC-Debate+ has the highest average improvement compared to all other methods. Additionally, the improvement gained by methods such as SoM, SFT or DebateGPT is far less stable than that of ACC-Debate+. In particular, for Mistral, SoM yields nearly no improvement, similarly SFT and DebateGPT offer little improvement when applied to of Llama-3.
Figure 3: Accuracy over five rounds of debate on BoolQ.
@@
Next, we look more closely at the accuracy of each method across five rounds of debate. Figure 3, shows per-round accuracy on the BoolQ dataset. As already illustrated by Table 1, ACC-Debate and ACC-Debate+ achieve higher final round accuracy than the other methods. Notably, our method has higher accuracy both at the final round $t=4$ and at round $t=0$. Recall that at round $t=0$, the actor’s response is independent of the critic. This indicates that in some cases, our training pipeline can produce actor models with superior zero-shot accuracy (without debate) compared to models produced by SFT and DebateGPT.
Interestingly, we observe that in some cases, SFT and DebateGPT are not much better than simple debate with untrained models (i.e., SoM-4x), e.g., BoolQ with Llama-3. This is primarily due to the fact that these methods are designed to improve the model’s single-short performance but do little to improve their collaborative abilities. Cases such as this outline the necessity of training to improve collaboration (ACC-Debate) rather than training for raw accuracy (SFT and DebateGPT).

### v3

@@ 4.4 Guided Collaborative Trajectories
To address these limitations and improve efficiency, we propose Guided-Collaborative Trajectories, which steer the deliberation procedure in two opposing directions: one towards, and another away from, the correct. By comparing these guided trajectories with the natural deliberation trajectory, we can assess the relative goodness of each trajectory using an estimation of the reward structure $r$.
@@
Specifically, for task $x$ with answer $y$, let $\mathbf{z}^{(t-1)}=(z_{a}^{(t-1)},z_{c}^{(t-1)})$ be the agents’ responses at time $t-1$. Let $(z_{a}^{(t)},z_{c}^{(t)})$ be the agents’ natural responses (i.e., without guidance), let $(z_{y,a}^{(t)},z_{y,c}^{(t)})$ and $(z_{!y,a}^{(t)},z_{!y,c}^{(t)})$ be the agents responses when guided towards, and away from, supporting answer $y$ respectively. Thus, each guided response is an off-policy generation. In practice, we want guided responses to be different enough from natural responses so that learning the guided responses results in consequential changes to the agent, but not so different that they are challenging to learn; we find that prompt modification is an effective tool for striking this balance. To guide the generations of $(z_{y,a}^{(t)},z_{y,c}^{(t)})$ and $(z_{!y,a}^{(t)},z_{!y,c}^{(t)})$, we will simply provide a correct and wrong target answer in the prompt, respectively - see “Guided Collaborative Trajectory Prompts” in Section C for further details.
@@
The terms $\Delta_{y}$ and $\Delta_{!y}$ give the expected accuracy difference if at round $t$ the actor had given response $z_{y,a}^{(t)}$ (or $z_{!y,a}^{(t)}$) instead of response $z_{a}^{(t)}$. Large $\Delta_{y}$ indicates that a one-response difference during the deliberation was sufficient to push the procedure toward the correct answer. Such responses would be desirable for the agent to learn. On the other hand, large values of $\Delta_{!y}$ indicate that a one-response difference easily causes the agents to converge to the incorrect answer; this indicates that the deliberation procedure is particularly fragile at timestep $t$.
With these observations in hand, we use $\Delta_{y}$ and $\Delta_{!y}$ to define positive an negative examples, in particular for a threshold $\varepsilon$,
$(z_{+}^{(t)},z_{-}^{(t)})=\begin{cases}(z_{y}^{(t)},z^{(t)})&\text{if}\quad\varepsilon\leq\Delta_{y}=r(z_{y,a}^{(t)},x,y)-r(z_{a}^{(t)},x,y)\\ (z^{(t)},z_{!y}^{(t)})&\text{if}\quad\varepsilon\leq\Delta_{!y}=r(z_{a}^{(t)},x,y)-r(z_{!y,a}^{(t)},x,y)\end{cases}$ (5)
if neither value is above the threshold, then the example is thrown out.
@@
Under Eq. 5, a positive example $z_{+}^{(t)}$ can be interpreted as a guided response $z_{y}^{(t)}$ which increased the probability of deliberation converging to the correct answer by at least $\varepsilon$, when compared with the natural response $z^{(t)}$. Similarly, a negative example $z_{-}^{(t)}$ is a guided response $z_{y!}^{(t)}$ which decreased the probability of deliberation converging to the answer by at least $\varepsilon$.
Now that we have a procedure for generating high-quality training examples consisting of positive and negative pairs, we next discuss how to use those positive and negative pairs to train both agents.
4.5 Learning on Guided Trajectories
@@
Hence, given a preference dataset of positive and negative examples for both the actor and critic agent, of the from $z_{-}^{(t)},z_{+}^{(t)}$, the DPO loss is defined as,
$\mathcal{L}_{\text{DPO}}=\sum_{t=0}^{T}\mathbb{E}_{(x,y,z_{-}^{(t)},z_{+}^{(t)})\sim D}\bigg{[}\log\sigma\bigg{(}\frac{\pi_{\theta}\big{(}z_{+}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}{\pi_{\text{ref}}\big{(}z_{+}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}-\frac{\pi_{\theta}\big{(}z_{-}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}{\pi_{\text{ref}}\big{(}z_{-}^{(t)}|x,\mathbf{z}^{(t-1)}\big{)}}\bigg{)}\bigg{]}$ (6)
Where $\pi_{\theta}$ is the policy induced by parameters $\theta_{a}$ or $\theta_{c}$ and $\mathbf{z}^{(t-1)}$ are the agent’s responses at the previous round (i.e., the responses prior to giving either response $z_{-}^{(t)}$ or $z_{+}^{(t)}$).
@@
Recall, given our generated preference data, this loss implicitly optimizes the reward $r(z^{t},x,y)$ which itself is equivalent to final accuracy (i.e., the quantity being maximized by Eq. 1). By summing across all rounds, we implicitly maximize the probability that each round $t$ yields a response $z^{t}$ which causes deliberation to converge to the correct answer at time $T$.
@@ 5 Experiments
[…]
@@
We compare ACC-Collab to several multi-agent and single-agent baselines. For inference-based methods, we compare to Society of Minds SoM Du et al. (2023), Persona Chan et al. (2023). For training-based methods, we compare to supervised fine-tuning SFT Radford (2018), DebateTune Li et al. (2024b), DebateGPT Subramaniam et al. (2024). We use three different base models: Llama-3-8B-Instruct Llama-3 Dubey et al. (2024), Mistral-7B-Instruct Mistral Jiang et al. (2023), and Gemma-2-2B-Instruct Gemma-2 Team et al. (2024).
5.1 ACC-Collab Performance
Final Answer Accuracy
@@ 5.1 ACC-Collab Performance
We begin by examining the performance of our method ACC-Collab (a single round of training) and ACC-Collab+ (two rounds of training) In table 1, we see the average accuracy of each method after five rounds of deliberation. Our method attains superior performance compared with baseline methods in most cases. The high efficacy of ACC-Collab relative to the baselines indicates that in most cases, only a single round of training is necessary to produce a high-quality collaborative team. It is worth noting that in some cases, further training rounds may decrease performance (i.e., ACC-Collab+ can have worse performance than ACC-Collab). Hence, we have a hold-out set of tasks to determine whether further training degrades performance.
Process Accuracy
In addition to measuring the correctness of the actor’s final answer (i.e., outcome accuracy), we also analyze the correctness of the team’s reasoning and discussion steps (i.e., process accuracy). Since ground truth is not available for reasoning and discussion steps, we use GPT-4o as an oracle to evaluate whether the agents follow the correct steps to reach the final answer. We provide an outline of the experimental setup and full results in Section A.1 of the supplement. We find that our method improves or maintains the process accuracy of the actor and critic.
@@
Llama-3             (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Collab ACC-Collab+ BoolQ .$812_{\pm.01}.$$811_{\pm.007}.$$781_{\pm.002}.$$775_{\pm.033}.$$798_{\pm.006}.$$815_{\pm.005}.$$887_{\pm.005}.$$\mathbf{.894}_{\pm.003}$ MMLU .$62_{\pm.004}.$$635_{\pm.004}.$$639_{\pm.004}.$$63_{\pm.004}.$$642_{\pm.005}.$$654_{\pm.005}.$$644_{\pm.01}.$$\mathbf{.683}_{\pm.012}$ BBH .$508_{\pm.003}.$$514_{\pm.005}.$$509_{\pm.013}.$$508_{\pm.005}.$$552_{\pm.006}.$$551_{\pm.008}.$$\mathbf{.593}_{\pm.006}$$574_{\pm.003}.$ SCIQ .$925_{\pm.002}.$$923_{\pm.002}.$$925_{\pm.004}.$$924_{\pm.004}.$$925_{\pm.003}.$$932_{\pm.001}.$$\mathbf{.952}_{\pm.0}$$948_{\pm.003}.$ ARC .$874_{\pm.001}.$$874_{\pm.001}.$$87_{\pm.003}.$$871_{\pm.002}.$$879_{\pm.004}.$$876_{\pm.002}.$$\mathbf{.881}_{\pm.004}$$869_{\pm.002}.$ Mistral               (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Collab ACC-Collab+ BoolQ .$801_{\pm.005}.$$798_{\pm.004}.$$831_{\pm.003}.$$83_{\pm.003}.$$84_{\pm.003}.$$848_{\pm.002}.$$877_{\pm.002}.$$\mathbf{.893}_{\pm.002}$ MMLU .$57_{\pm.003}.$$562_{\pm.005}.$$574_{\pm.002}.$$562_{\pm.005}.$$594_{\pm.004}.$$577_{\pm.002}.$$61_{\pm.005}.$$\mathbf{.672}_{\pm.004}$ BBH .$428_{\pm.002}.$$462_{\pm.003}.$$465_{\pm.011}.$$456_{\pm.005}.$$439_{\pm.006}.$$48_{\pm.012}.$$519_{\pm.009}.$$\mathbf{.601}_{\pm.004}$ SCIQ .$856_{\pm.002}.$$856_{\pm.002}.$$86_{\pm.002}.$$863_{\pm.003}.$$858_{\pm.004}.$$871_{\pm.003}.$$902_{\pm.005}.$$\mathbf{.905}_{\pm.002}$ ARC .$824_{\pm.001}.$$823_{\pm.002}.$$827_{\pm.001}.$$834_{\pm.0}.$$825_{\pm.003}.$$822_{\pm.002}.$$843_{\pm.003}.$$\mathbf{.856}_{\pm.003}$ Gemma-2              (Ours)      (Ours) SoM (2x) SoM (4x) Persona DebateTune SFT DebateGPT ACC-Collab ACC-Collab+ BoolQ .$75_{\pm.011}.$$759_{\pm.004}.$$716_{\pm.015}.$$767_{\pm.003}.$$783_{\pm.011}.$$812_{\pm.003}.$$84_{\pm.005}.$$\mathbf{.845}_{\pm.005}$ MMLU .$58_{\pm.002}.$$578_{\pm.002}.$$577_{\pm.002}.$$578_{\pm.001}.$$579_{\pm.002}.$$\mathbf{.582}_{\pm.002}$$51_{\pm.016}.$$555_{\pm.003}.$ BBH .$454_{\pm.007}.$$449_{\pm.01}.$$447_{\pm.006}.$$447_{\pm.007}.$$498_{\pm.006}.$$491_{\pm.01}.$$\mathbf{.513}_{\pm.006}$$475_{\pm.008}.$ SCIQ .$903_{\pm.002}.$$903_{\pm.002}.$$908_{\pm.003}.$$903_{\pm.001}.$$913_{\pm.002}.$$914_{\pm.002}.$$\mathbf{.918}_{\pm.003}$$909_{\pm.003}.$ ARC .$841_{\pm.003}.$$843_{\pm.005}.$$847_{\pm.003}.$$847_{\pm.003}.$$848_{\pm.002}.$$851_{\pm.003}.$$\mathbf{.852}_{\pm.003}$$849_{\pm.002}.$
@@
Table 1: Average accuracy (with 95% confidence intervals) after $5$ rounds of deliberation. For each dataset, the highest accuracy is shown in bold.
5.2 Performance Increase of Multi-Agent Collaboration
@@
Figure 2: Percent improvement in accuracy after five rounds of deliberation, compared to a single round. Percent improvement (Eq. 7) for each method is averaged across all five datasets.
Average Improvement
@@
As noted in Du et al. (2023), the key mechanism behind the success of multi-agent deliberation (or any of its many variants) is that discussion over multiple rounds allows the models to iteratively refine their answers. Thus, a natural question for any iterative multi-agent method is: how much does accuracy improve from the initial round $t=0$ to the final round $t=T$? Where $T=4$ in our experiments. To measure this, we look at the percent improvement in model accuracy from round $t=0$ to round $t=4$ calculated as,
@@
In Figure 2 we see the average percent improvement for each method, averaged across all 5 datasets. For each of the three base models, ACC-Collab+ has the highest average improvement compared to all other methods. Additionally, the improvement gained by methods such as SoM, SFT or DebateGPT is far less stable than that of ACC-Collab+. In particular, for Mistral, SoM yields nearly no improvement, similarly SFT and DebateGPT offer little improvement when applied to of Llama-3.
Figure 3: Accuracy over five rounds of deliberation on BoolQ (top) and SCIQ (bottom).
@@
Next, we look more closely at the accuracy of each method across five rounds of deliberation. Figure 3, shows per-round accuracy on the BoolQ dataset. As already illustrated by Table 1, ACC-Collab and ACC-Collab+ achieve higher final round accuracy than the other methods. Notably, our method has higher accuracy both at the final round $t=4$ and at round $t=0$. Recall that at round $t=0$, the actor’s response is independent of the critic. This indicates that in some cases, our training pipeline can produce actor agents with superior zero-shot accuracy (without deliberation) compared to models produced by SFT and DebateGPT.
Interestingly, we observe that in some cases, SFT and DebateGPT are not much better than simple deliberation with untrained models (i.e., SoM-4x), e.g., BoolQ with Llama-3. This is primarily due to the fact that these methods are designed to improve the model’s single-short performance but do little to improve their collaborative abilities. Cases such as this outline the necessity of training to improve collaboration (ACC-Collab) rather than training for raw accuracy (SFT and DebateGPT).